OpenStudy (anonymous):

what equation do use to find the line that passes thur (-2,-1) and (-4,-3)

OpenStudy (anonymous):

y - y1 = m(x - x1)

OpenStudy (anonymous):

That's point-slope form. First, Do you remember the formula for finding m, the slope?

OpenStudy (anonymous):

When considering which formula you should use, always look at what you are given first....then choose.

OpenStudy (anonymous):

m=y2-y1/x2-x1